OpenStudy (anonymous):

prove cos(3pi/2 + x) = sin(x)

OpenStudy (anonymous):

Since...it is a Odd multiple...Cos will Change to Sin and Since it is in Fourth Quadrant Cos is +ve so cos(3pi/2 + x) = Sin x

OpenStudy (anonymous):

cos 3 pi /2 = 0 sin 0 = 0 therefore cos 3 pi/2 = sin 0 now add x to both sides cos ( 3 pi/2 + x) = sin x

hero (hero):

You can't add x to both sides like that

hero (hero):

That's not a proof by the way

OpenStudy (anonymous):

x is variable if we add constants then why not variables

OpenStudy (anonymous):

use the sum or difference formula this is trig

hero (hero):

You cannot add x to both sides in the manner that you have @uzumakhi. Either way your steps don't really prove much.

OpenStudy (anonymous):

can you tell me a proper derivation of this? @Hero

hero (hero):

Actually, upon further review, it looks like your steps might be more legit than previously thought.

hero (hero):

Although when setting two things equal to zero, things get kinda tricky and fuzzy.

OpenStudy (anonymous):

is it cos(3pi/2 + x) = cos 3pi/2cosx - sin3pi/2sinx?

OpenStudy (anonymous):

if we see graphically then both cos 3pi/2 and sin 0 have same value that's why i do this

hero (hero):

I'm gonna have to check with an expert on this to see if those steps are valid for a proof.

hero (hero):

Okay, I have the reason why you cannot add x inside a function like that

hero (hero):

to both sides

OpenStudy (anonymous):

so tell me

hero (hero):

\[(1)^2 = (-1)^2\] is true However, (1 + x)^2 \ne (-1 + x)^2\]

hero (hero):

\[(1 + x)^2 \ne (-1 + x)^2\]

hero (hero):

actually, wait

hero (hero):

\[-1 \ne 1\] to begin with The expert pointed that out, but in this case \[\cos(3pi/2) = \sin(0)\] so you're still winning in this regard.

hero (hero):

I think that cos(3pi/2) + x = sin(0) + x is valid but what you've done may not be

hero (hero):

But it still isn't really clear why your step is truly invalid

OpenStudy (anonymous):

but there is alternative derivation to prove this which should be appropriate

hero (hero):

Okay, here's a better and more convincing disproof of your methods

hero (hero):

sin(0) = 0 tan(0) = 0 sin(0) = tan(0) Add x to both sides sin(x) = tan(x)

OpenStudy (anonymous):

but there is also a case that sin and cos are mostly same